jim is 3 years older than his sister kim theirs ages add up to 13
how old is jim and how old is kim

Answer by stanbon(75887)

jim is 3 years older than his sister kim; their ages add up to 13
-------------------
Equations:
j + k = 13
j = k +3
-----
Substitute for "j" and solve for "k":
k+3 + k = 13
2k = 10
k = 5 yrs old (Kim's age now)
-----
Since j = k+3, Jim's age = 8 yrs
===================
Cheers,
Stan H.